Patient BN is a 36-year-old female with a type of organ failure that reduces her quality of life to half of what it would be in good health. Without treatment she can expect to live only two years. With a successful transplant, BN can expect to live four years and have a quality of life that is near 80% what she would enjoy in good health. However, the transplant costs $100,000, plus $10,000 each year for drugs and follow-up care, and carries a 15% risk of rejection resulting in immediate death.
a) What is the cost per additional year of life gained (without discounting for time or quality of life)?
b) What is the cost per discounted Quality-adjusted Life-year (QALY) gained (assuming a 5% time discount rate)?
Answer #1

a. Cost per additional year life gained without discounting for time and quality of life

Transplant cost= $100000

Annual medical costs per year=$10000

Cost per year of living upto 4 years = $140000/4 = $35000

b. Cost per QLAY, at 5% time discount rate, for 4 years

= 4*(0.8/(1+0.04)) = 3.076

calculate cost at the actual discounted rate of 3.076

$140000/3.076 = $45513 is the cost per discounted QLAY gained.